OpenStudy (anonymous):

Is there a thing as an inverse Log? I mean like Sinx=y, then inverse is Sin^(-1)y=x is there same thing for log?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

ln?

OpenStudy (zarkon):

\[a^{\log_a(x)}=x\] \[\log_{a}(a^x)=x\]

OpenStudy (anonymous):

So if I have \[\log_3(\log_x)=5\] How would I do that?

OpenStudy (zarkon):

what is inside the inner log?

OpenStudy (anonymous):

O, forgot to put it, 3

OpenStudy (zarkon):

log base 3 of x?

OpenStudy (anonymous):

\[\log_3(\log_x3)=5\]

OpenStudy (zarkon):

ok... \[3^{\log_{3}(\log_{x}(3))}=\log_{x}(3)\]

OpenStudy (zarkon):

that gives \[\log_{x}(3)=3^5\]

OpenStudy (anonymous):

how did you get your first thing, 3 to the log base 3.... ?

OpenStudy (zarkon):

\[\log_3(\log_x3)=5\Rightarrow 3^{\log_3(\log_x3)}=3^5\]

OpenStudy (anonymous):

I don't get that, sorry.

OpenStudy (zarkon):

if \[a=b\] then \[3^a=3^b\]

OpenStudy (anonymous):

Ok, I get it, and what's after you got that second equation.

OpenStudy (zarkon):

\[\log_3(\log_x3)=5\Rightarrow 3^{\log_3(\log_x3)}=3^5\] \[\Rightarrow \log_{x}(3)=3^5\]

OpenStudy (zarkon):

Then use \[\log_a(b)=c \Leftrightarrow a^c=b\]

OpenStudy (anonymous):

x^5=3^5 x=3

OpenStudy (anonymous):

Oh my bad

OpenStudy (zarkon):

\[\Rightarrow \log_{x}(3)=3^5\Leftrightarrow x^{(3^5)}=3\]

OpenStudy (anonymous):

OK so x=3/125

OpenStudy (zarkon):

so \[x^{243}=3\]

OpenStudy (anonymous):

Oh the other way, yes 81

OpenStudy (zarkon):

no

OpenStudy (zarkon):

raise both sides to the 1/243 power

OpenStudy (anonymous):

243th root of 3

OpenStudy (anonymous):

Should have got that before....

OpenStudy (zarkon):

yes\[\sqrt[243]{3}\]

OpenStudy (anonymous):

Yep. I meant that, ty!